# On the gravitational stratification of multi-fluid-multi-species plasma

### v1

@@ 3.1 Revisiting the SE assumption
We calculate an SE stratification by first extracting the temperature distribution from the widely used Model C7 of Avrett & Loeser (2008). The total number density at the bottom of the photosphere ($z=0$ Mm) is approximated using the hydrogen number density given by Model C7, as ion and electron number densities are still small on the solar surface. In addition to the four fluids included in Section 2, we further include Ne, Ne+, Fe, and Fe+. Then, SE ionization under the given temperature distribution is calculated to provide the number densities of these eight fluids, using the ionization and recombination rates of Voronov (1997) and McWhirter (1965), with photospheric abundances of Asplund et al. (2021) provided through the CHIANTI atomic database, Version 11.0 (Dufresne et al., 2024). Note that we calculate only the cHE distributions here, because pHE would lead to significant deviations in the resulting ionization fractions, as explained in the previous section. The cHE stratification is shown in Figure 2.
We find that the total ion density below the transition region (TR) is around 10 orders of magnitude lower than the total neutral density, which relates to several factors. Firstly, the ionization state of hydrogen and helium is computed in SE including collisional ionization balanced by radiative and three body recombinations (for heavier ions dielectronic recombinations would also be included), using rates from the CHIANTI package (Del Zanna & Mason, 2018), while leaving out photoionizations. In addition, as mentioned, SE may underestimate ion fractions due to slow ionzation and recombination rates for some transitions. Thus, in the upper chromosphere, NEQ ionization is required, while in the photosphere, solving the Saha equation (Saha, 1920) including all species is needed, as metals with their low FIP are the main donors of electrons. A more realistic calculation of SE ionization was presented by Gómez-Míguez et al. (2024), who used the MULTI code (Carlsson, 1986) to solve the radiative transfer equation together with the rate equations thus including photoionization correctly. However, calculating the ionization fractions is not the goal of this work, and Figure 2 shows that, most importantly, over the whole domain, at each height, the ionization fractions are in SE. Therefore, the present cHE stratification is very close to both hydrostatic equilibrium and SE. In the following, we further show that a cHE stratification can be constructed for more complicated settings.
@@
Figure 3: A snapshot of a rMHD simulation with NEQ ionization. Top: density; middle: temperature; bottom: magnetic field strength. Loops and spicules are observed in the snapshot. These structures are dynamic but also relatively steady compared to, e.g., high-frequency waves that may be responsible for local heating or the FIP effect.
@@
The previous stratification is given directly using the ionization and recombination rates in SE, but the ionization fractions of different species can be extracted from, e.g., Bifrost simulations. Here we prove the flexibility of the present numerical integration routine, which is applicable as long as temperature and ionization fraction distributions are provided.
@@
We adopt a NEQ simulation snapshot from the 2.5D rMHD model Martínez-Sykora et al. (2020a), as shown in Figure 3. In this model, we observe $\sim$50 Mm long loops, which are associated with the underlying magnetic field configuration, which has two main opposite polarities. Correspondingly, there are cold and dense features with chromospheric temperature along the loops, resembling spicules and jets (Martínez-Sykora et al., 2017). More details of the rMHD simulation are, however, not further discussed here. We calculate the horizontally-averaged density and temperature, and then again use the method introduced in Section 3.1 to obtain the SE ionization fractions. Note that the result here includes two-times ionized helium (He2+) but not iron or neon, resulting in a two-species-five-fluid mixture. For comparison, we also extract the averaged ionization fractions directly from the 2.5D snapshot, replacing the SE ionization fractions, to perform the same integration routine. We obtain two stratifications shown in Figure 4, while omitting the contribution of magnetic pressure gradient. The inclusion of magnetic pressure is briefly discussed in Appendix A. Note that both stratifications are calculated using the same temperature distribution.
@@
Figure 4: Two cHE stratifications calculated based on the rMHD simulation snapshot. Left: ionization fractions are calculated assuming SE and using the averaged total density and temperature from the rMHD model; right: NEQ ionization fractions are directly given as the averaged ionization fractions from the rMHD model. A small value is used as the floor when calculating ionization fractions in SE, which is a practical treatment for simulations but serves here only for the convenience of visualization, without affecting our conclusions.
@@
Due to the horizontal averaging, we obtain a relatively flat TR from the rMHD model. Still, the aim is to show that a cHE stratification can be obtained to include effects of NEQ ionization. Specifically, the NEQ stratification has higher ion populations in the chromosphere and photosphere. Note that the high ion fractions in the NEQ result are related to the long time-scales of the ionization and recombination processes, as well as their interplay with chromospheric dynamics, e.g., shocks (Carlsson & Stein, 1992, 2002; Leenaarts et al., 2007; Wedemeyer-Böhm & Carlsson, 2011). Therefore, the present numerical integration routine provides flexibility for specifically studying atmospheric dynamics while considering the NEQ ionization effects. In contrast, a pHE stratification cannot recover such effects, as shown in Figure 1.
@@ 4 MFMS equations and numerical methods
In the following, we run 1D MFMS numerical simulations that use the stratifications considered above as initial conditions. We use the numerical code Ebysus (Martínez-Sykora et al., 2020b, 2023; Wargnier et al., 2023, 2025). Compared to various multi-fluid codes (Alvarez Laguna et al., 2016; Popescu Braileanu et al., 2019; Popescu Braileanu & Keppens, 2022; Snow et al., 2023), an important feature of Ebysus is that it can solve the MHD equations separately for each of the desired number of excited levels, ionization stages and species, although in this work, we solve the collective motion of all the excited levels of each neutral species for simplicity. Ebysus is modular and allows for the inclusion of ionization/excitation in NEQ, collision between different fluids, thermal conduction, radiative losses, and the Hall term. Details of the derivation of the MFMS formulae have been introduced by Khomenko et al. (2014); Ballester et al. (2018). For the purpose of this work, we neglect the Hall term and thermal conduction as well as considering only 1D as detailed in the next two sections.
@@
where $q_{a{I}}$ is the ion charge. Note that, while considering fewer species, similar formulae were also used by Martínez-Gómez et al. (2016).
More importantly, Ebysus solves a separated energy equation of electrons along with the energy equations of ions and neutral fluids. Consequently, the contribution of electron pressure (gradient) is not directly included in the momentum and energy equations of ions. Whereas, the electron pressure contributes to the electric field, which is then included in the momentum equations of ions. Thus, electron pressure also contributes to the induction equation, which reads
@@ 5.1 Initial conditions
A cHE stratification obtained using the numerical integration routine is approximately in equilibrium. Therefore, there will be remnant disturbances when a simulation uses the cHE stratification as the initial condition. More importantly, according to our assumption, there must be drift velocities between fluids, which are related to the total pressure gradient in the model. To evaluate these effects, we simulate disturbances developing in initially quiet stratifications, without imposing any external dynamic driving force. Specifically, we construct a cHE stratification using the Model C7 temperature distribution, yet including only [^3] The Model C7 provides discrete data points. Therefore, the temperature distribution is interpolated using the interp1d function from the NumPy library (Harris et al., 2020) with the quadratic interpolation method. While errors still exist, using the linear method would have caused more significant errors in drift velocities and, eventually, in the total velocity. The error in the interpolated temperature distribution is not observable by the naked eye, but it affects the sound speed and pressure gradient. 2.2 to 3.2 Mm and hydrogen and helium (no He2+). Therefore, the domain does not include the non-smooth transition from the upper chromosphere to the TR at around 2.14 Mm in the Model C7. We include only the smooth (yet steep) variation above the transition, where pressure gradient is large.
Figure 5: Left: density distributions at $t=0$ s and $t=0.5$ s when using the cHE stratification as the initial condition for a simulation with NEQ ionization/recombination, but without any dynamic driving force; right: the pHE stratification (viz. at $t=0$ s). Using the pHE stratification as the initial condition leads to numerical instability in simulations with ionization/recombination, and thus the results are not included.
@@
The resulting initial cHE stratification (Figure 5, left panel in solid lines) is similar to the result shown in Figure 2 for the selected domain. The corresponding simulation result at $0.5$ s (Figure 5, left panel in dashed lines) is also shown, and is discussed later. Across the whole domain of the cHE stratification, plasma is approximately in SE, and thus higher up in the domain the helium ionization fraction continues to increase with increasing temperature. Correspondingly, the fractions of neutrals gradually decrease. For comparison, the initial pHE stratification constructed using the same plasma conditions at $z=2.2$ Mm as the cHE stratification is shown with solid lines in Figure 5, right panel. In the initial pHE stratification, the ionization fractions do not change with the (rapidly) increasing temperature, and thus the plasma is not in SE.
@@ 5.2 SE evolution
Using the given initial conditions we run simulations without any additional dynamic driving force. The simulations use physically calculated collisional frequencies (Wargnier et al., 2022), and include the NEQ ionization and recombination processes. We compare the temporal variations of densities in the simulation using the cHE stratification as the initial condition, as shown with dashed lines in Figure 5, left panel. The density distributions change little over the simulated time, except near the lower boundary, where more accurate spatial discretization is usually needed (Krause, 2019; Wei et al., 2023). Therefore, we may claim that the initial condition is sufficiently close to SE in the whole domain. However, the simulation starting with the pHE stratification develops numerical instability as a result of the rapid ionization caused by the initial nonequilibrium, and thus the corresponding result is not further discussed.
Although the total velocity in the cHE + NEQ simulation is not exactly zero, only small disturbances slowly develop over time, mainly in the lower part of the domain (Figure 6). The decoupling in the cHE result causes collisional forces. In the meantime, the ionization and recombination processes exchange momentum between ions and neutrals, reducing the decoupling. Without ionization and recombination, the decoupling increases over time at a faster rate, until the collisional forces are able to balance the gravity and the pressure gradient, in which case the maximum decoupling between the neutral helium and the total velocity reaches above 2 km s-1 (not shown). The drift velocities that occur in the cHE stratification are physically important. In particular, such drift velocities are an important difference between single-fluid and multi-fluid models (Gómez-Míguez et al., 2025).
The results suggest that the pHE stratification is not viable for simulations that include ionization and recombination, as numerical instability may occur. Even if ionization and recombination are not activated, the ionization fractions are important for correctly describing the collisional interactions. For example, in Figure 5, the density of the neutral hydrogen in the pHE stratification can be around an order of magnitude higher that of the neutral hydrogen in the cHE stratification, and this difference certainly affects collisional frequencies.
@@ 6 The ponderomotive force and the resulting motions
Since significant decoupling appears in the cHE stratification when the pressure gradient is large, it is interesting to know how this ”static” decoupling would contribute to the MFMS dynamics. Here, we run simulations of a monochromatic Alfvén wave, using the cHE and pHE stratifications introduced in the previous section. Alfvén waves existing in the low solar atmosphere have been suggested (De Pontieu et al., 2007; McIntosh et al., 2011; Chae et al., 2021), and they may play a vital role in coronal heating (Van Doorsselaere et al., 2020), the FIP effect (Laming, 2004, 2015), etc., which are beyond the scope of this work. We are interested in the fact that the ponderomotive force caused by Alfvén waves drives plasma motions along the magnetic field, and such motions can be simulated and used as an example to compare the effects of using different stratifications.
An initially constant vertical magnetic field $B_{z}=10$ G is imposed on the stratifications discussed in the previous section, and the Alfvén wave is driven by periodically changing $B_{x}$ at $z=2.2$ Mm, similar to what was done by (Martínez-Sykora et al., 2023). Note that such a driver may not be suitable in the weakly ionized photosphere (Cally, 2023), but works under the present atmospheric condition. The frequency of the driver is 1 Hz, and the amplitude is 1 G. The resulting amplitude of the velocity along the $x$ direction is $\sim 40$ km s-1. Based on this basic setting, five simulations are performed, as listed in Table 1. We mainly focus on comparing the differences between using cHE and pHE stratifications, with or without the NEQ ionization (simulations (c)-(e)). Because of the numerical instability mentioned in the previous section, the pHE stratification is only discussed without including ionization and recombination in the simulation (simulation (d)). We also include two simulations that use non-physical collisional frequencies to examine the weakly and strongly collisional scenarios (simulations (a) & (b)). These two non-physical frequencies are, respectively, significantly lower and higher than the imposed wave frequency. The former should evolve as separated fluids and the latter should recover the single fluid scenario.
@@
Figure 7: Longitudinal velocities at $t=1$ s. All settings are listed in Table 1. Except simulation (d) which uses the pHE stratification, all the others use the cHE stratification as the initial condition. Simulations (a) and (b) use assigned constant collisional frequencies. More specifically, (a) is weakly collisional and (b) is strongly collisional, respectively having collisional frequencies $5\times 10^{-2}$ Hz and $10^{2}$ Hz. All the other simulations use physical collisional frequencies calculated by Ebysus. Only simulation (e) includes the NEQ ionization/recombination.
@@
The effects of the numerical disturbances appear mainly in the $z$ direction (i.e., the direction of the magnetic field). Moreover, the Alfvén wave induces ponderomotive force, driving plasma motions along the magnetic field. Therefore, the longitudinal velocity components are compared in Figure 7. Apparently, when it is weakly collisional (panel a), the neutral fluids are decoupled from the ionized fluids, whose motions are driven by the Alfvén wave. Each neutral fluid also moves independently because of the imbalance between gravity and the respective plasma gas pressure, while neither collisions with ionized fluids nor magnetic pressure affect neutral fluids. In contrast, all fluids in the strongly collisional simulation are coupled, behaving as a single-fluid as expected (panel b). More importantly, we can observe that in the strongly collisional simulation, the total velocity is zero before the wave arrives, showing that the cHE stratification is indeed in hydrostatic equilibrium, providing another way to validate the MFMS code – once it is collisionally coupled the plasma behaves as a single fluids.
@@
Simulations using physical collisional frequencies (Wargnier et al., 2022) show significant decoupling of neutral helium (panels c-e). A similar result was also obtained by Martínez-Gómez et al. (2018), although only initially homogeneous media were considered. Wargnier et al. (2023) also found that neutral helium is more significantly decoupled from other fluids during upper-chromospheric magnetic reconnection. Thus, including helium may be important for modeling the dynamics across the TR. Moreover, when using cHE stratification (panels c and e), neutral helium has a higher longitudinal velocity than pHE (panel d), due to the contribution of the pressure gradient, which is more obvious in Figure 6. The ionization process strengthens the interactions between ionized and neutral fluids, thereby increasing the coupling in the simulation (e). Nevertheless, in general, neutral helium has smaller longitudinal velocity than the other fluids.
@@
The pHE stratification is not further discussed with ionization and recombination due to the numerical instability. We should however note that, the ionization and recombination timescales are long in the chromosphere (Carlsson & Stein, 2002), and thus the changes in density and temperature would be less drastic when the pHE stratification is used under chromospheric conditions. However, this also means that the ionization fractions (and thus the collisional interactions) in the corresponding simulations would be largely dependent on the initial stratification. We have shown in Section 2, that heavier elements, whose scale heights are different from that of (ionized or neutral) hydrogen, cannot be properly recovered by the pHE stratification, even if the collisional interactions (or other mixing processes) are sufficient to couple the heavy fluids.

### v2

@@ 3.2 Revisiting the SE assumption
We calculate an SE stratification by first extracting the temperature distribution from the widely used Model C7 (Avrett & Loeser, 2008). The total number density at the bottom of the photosphere ($z=0$ Mm) is approximated using the hydrogen number density given by Model C7, as ion and electron number densities are still small on the solar surface. In addition to the four fluids included in Section 2, we further include [^4] There is effectively no limitation on the number of species that Ebysus can include, as long as relevant information (e.g. collisional cross sections) are available. Different species are included in this work simply to show the flexibility of the present approach. Ne, Ne+, Fe, and Fe+. Then, SE ionization under the given temperature distribution is calculated to provide the number densities of these eight fluids, using the ionization and recombination rates of Voronov (1997) and McWhirter (1965), with photospheric abundances of Asplund et al. (2021) provided through the CHIANTI atomic database, Version 11.0 (Dufresne et al., 2024). Note that we calculate only the cHE distributions here, because pHE would lead to significant deviations in the resulting ionization fractions, as explained in the previous section. The cHE stratification is shown in Figure 2.
We find that the total ion density below the transition region (TR) is around 10 orders of magnitude lower than the total neutral density, which relates to several factors. Firstly, the ionization state of hydrogen and helium is computed in SE, including collisional ionization balanced by radiative and three-body recombinations (for heavier ions, dielectronic recombinations would also be included), using rates from the CHIANTI package, while leaving out photoionizations. In addition, as mentioned, SE may underestimate ion fractions due to slow ionization and recombination rates for some transitions. Thus, in the upper chromosphere, NEQ ionization is required, while in the photosphere, solving the Saha equation (Saha, 1920) including all species is needed, as metals with their low FIP (first ionization potential) are the main donors of electrons. A more realistic calculation of SE ionization was presented by Gómez-Míguez et al. (2024), who used the MULTI code (Carlsson, 1986) to solve the radiative transfer equation together with the rate equations, thus including photoionization correctly. However, how the ionization fractions are calculated does not affect the numerical integration. Figure 2 shows that, over the whole domain, the ionization fractions are in SE. Therefore, the present cHE stratification is in both hydrostatic equilibrium and SE.
@@
Figure 3: A snapshot of a rMHD simulation with NEQ ionization. Loops and spicules are observed in the snapshot. These structures are dynamic but also relatively steady compared to, e.g., high-frequency waves that may be responsible for local heating or the FIP effect.
@@
The previous stratification is calculated using ionization fractions based on the ionization and recombination rates in SE. However, NEQ ionization fractions can also be used to calculate the number densities needed for the numerical integration. Below we prove the flexibility of the present numerical integration routine, which is applicable as long as temperature and ionization fraction distributions are provided, irrespective of how they are calculated.
@@
We adopt a NEQ simulation snapshot from the 2.5D rMHD model (Martínez-Sykora et al., 2020a), as shown in Figure 3. In this model, we observe $\sim$50 Mm long loops, which are associated with the underlying magnetic field configuration, which has two main opposite polarities. Correspondingly, there are cold and dense features with chromospheric temperature along the loops, resembling spicules and jets (Martínez-Sykora et al., 2017). More details of the rMHD simulation are not further discussed here. We calculate the horizontally-averaged density and temperature, and then at first use the method introduced in Section 3.2 to obtain the SE ionization fractions. Note that the result here includes two-times ionized helium (He2+) but not iron or neon, resulting in a two-species-five-fluid mixture. Subsequently, for comparison, we also extract the averaged NEQ ionization fractions from the 2.5D snapshot, replacing the SE ionization fractions, to perform the same integration routine. We obtain two stratifications shown in Figure 4, while omitting the contribution of magnetic pressure gradient. The inclusion of magnetic pressure is briefly discussed in Appendix A. Note that both stratifications are calculated using the same temperature distribution.
@@
Figure 4: Two cHE stratifications calculated based on the rMHD simulation snapshot. Left: ionization fractions are calculated assuming SE and using the averaged total density and temperature from the rMHD model; right: NEQ ionization fractions are directly given as the averaged ionization fractions from the rMHD model. A small value is used as the floor when calculating ionization fractions in SE, which is a practical treatment for simulations, without affecting our conclusions.
@@
Due to the horizontal averaging, we obtain a relatively flat TR from the rMHD model. Still, the aim is to show that a cHE stratification can be obtained to include the effects of NEQ ionization. Specifically, the NEQ stratification has higher ion populations in the chromosphere and photosphere, resulting from the long time-scales of the ionization and recombination processes, as well as their interplay with chromospheric dynamics, e.g., shocks (Carlsson & Stein, 1992, 2002; Leenaarts et al., 2007; Golding et al., 2014; Wedemeyer-Böhm & Carlsson, 2011). Therefore, the present numerical integration routine provides flexibility for specifically studying atmospheric dynamics while considering the consequence of NEQ ionization. In contrast, a pHE stratification cannot recover such effects, as shown in Figure 1.
@@ 4 MFMS equations and numerical methods
In the following, we run 1D MFMS numerical simulations that use the stratifications considered above as initial conditions. We use the numerical code Ebysus (Martínez-Sykora et al., 2020b, 2023; Wargnier et al., 2023, 2025). Compared to various multi-fluid codes (Alvarez Laguna et al., 2016; Popescu Braileanu et al., 2019; Popescu Braileanu & Keppens, 2022; Snow et al., 2023), an important feature of Ebysus is that it can solve the MHD equations separately for each of the desired number of excited levels, ionization stages and species, although in this work, we solve the collective motion of all the excited levels of each neutral species for simplicity. Ebysus is modular and allows for the inclusion of ionization/excitation in NEQ, collision between different fluids, thermal conduction, radiative losses, and the Hall term. Details of the derivation of the MFMS formulae have been introduced by Khomenko et al. (2014); Ballester et al. (2018). For the purpose of this work, we neglect the Hall term and thermal conduction, and consider only 1D.
@@
where $q_{a{I}}$ is the ion charge. Note that, while considering fewer species, similar formulae were used by Martínez-Gómez et al. (2016).
More importantly, Ebysus solves a separated energy equation of electrons along with the energy equations of ions and neutral fluids. Consequently, the contribution of electron pressure (gradient) is not directly included in the momentum and energy equations of ions. Whereas the electron pressure contributes to the electric field, which is then included in the momentum equations of ions. Thus, electron pressure also contributes to the induction equation, which reads
@@ 5.1 Initial conditions
A cHE stratification obtained using the numerical integration routine is approximately in equilibrium. Therefore, there may be remnant disturbances when a simulation uses the cHE stratification as the initial condition. More importantly, according to our assumption, there must be drift velocities between fluids, which are related to the pressure gradients in the model. To evaluate these effects, we simulate disturbances developing in initially quiet stratifications, without imposing any external dynamic driving force. Specifically, we construct a cHE stratification using the Model C7 temperature distribution, yet including only [^5] The Model C7 provides discrete data points. Therefore, the temperature distribution is interpolated using the interp1d function from the NumPy library (Harris et al., 2020) with the quadratic interpolation method. While errors still exist, using the linear method would have caused more significant errors in drift velocities and, eventually, in the total velocity. The error in the interpolated temperature distribution is not observable by the naked eye, but it affects the sound speed and pressure gradient. 2.2 to 3.2 Mm and hydrogen and helium (no He2+). Therefore, the domain does not include the non-smooth transition from the upper chromosphere to the TR at around 2.14 Mm in the Model C7. We include only the smooth (yet steep) variation above the transition, where the pressure gradient is large.
Figure 5: Left: density distributions at $t=0$ s and $t=0.5$ s when using the cHE stratification as the initial condition for a simulation with NEQ ionization/recombination, but without any dynamic driving force; right: the pHE stratification (viz. at $t=0$ s). Using the pHE stratification as the initial condition leads to numerical instability in the simulation with ionization/recombination, and thus the result is not included.
@@
Figure 7: Temporal variations of the velocity distributions of neutral helium (solid lines) and ionized helium (dashed lines) in the cHE + NEQ simulation.
The resulting initial cHE stratification (Figure 5, left panel in solid lines) is similar to the result shown in Figure 2 for the selected range. Across the whole domain of the cHE stratification, plasma is approximately in SE, and thus, higher up in the domain, the helium ionization fraction continues to increase with increasing temperature. Correspondingly, the fractions of neutrals gradually decrease. For comparison, the initial pHE stratification constructed using the same plasma conditions at $z=2.2$ Mm as the cHE stratification is shown with solid lines in Figure 5, right panel. In the initial pHE stratification, the ionization fractions do not change with the (rapidly) increasing temperature, and thus the plasma is not in SE.
@@ 5.2 SE evolution
Using the given initial conditions, we run simulations without any additional dynamic driving force. The simulations use physically calculated collisional frequencies (Wargnier et al., 2022), and include the NEQ ionization and recombination processes. We compare the temporal variations of densities in the simulation using the cHE stratification as the initial condition, as shown with dashed lines in Figure 5, left panel. The density distributions change little over the simulated time, except near the lower boundary, where more accurate spatial discretization is usually needed (Krause, 2019; Wei et al., 2023). Therefore, we may claim that the initial condition is sufficiently close to SE in the whole domain. However, the simulation starting with the pHE stratification develops numerical instability as a result of the rapid ionization caused by the initial nonequilibrium, and thus the corresponding result is not further discussed.
Although the total velocity in the cHE + NEQ simulation is not exactly zero, only small disturbances slowly develop over time, more significantly in the lower part of the domain (Figure 6). The decoupling in the cHE result causes collisional forces. In the meantime, the ionization and recombination processes exchange momentum between ions and neutrals, reducing the decoupling. The velocity drift of neutral helium is the most significant, but its increase is slowing down over time, as shown in Figure 7. Particularly in the lower part of the domain, the velocities soon ($\sim$1 s) stop changing, and in the upper part, the relative drifts between ionized and neutral fluids also become stable. In the upper part, while both neutral and ionized helium move downwards, hydrogen moves upwards (not shown), and thus the total velocity remains around zero (Figure 6). In this specific case, the smaller pressure gradients higher up cause less decoupling, but the collisional interactions are also weaker to stabilize the decoupling. When without ionization and recombination, the decoupling changes over time at a faster rate, until the collisional forces are able to balance the gravity and the pressure gradients, in which case the maximum decoupling between the neutral helium and the total velocity reaches above 2 km s-1 (not shown). The drift velocities that occur in the cHE stratification are physically important. In particular, such drift velocities are an important difference between single-fluid and multi-fluid models (Gómez-Míguez et al., 2025).
The results suggest that the pHE stratification is not viable for simulations that include ionization and recombination, as numerical instability may occur. Even if ionization and recombination are not activated, the ionization fractions are important for correctly describing the collisional interactions. For example, in Figure 5, the density of the neutral hydrogen in the pHE stratification can be around an order of magnitude higher than that of the neutral hydrogen in the cHE stratification, and this difference certainly affects collisional frequencies.
@@ 6 The ponderomotive force and the resulting motions
Since significant decoupling appears in the cHE stratification when the pressure gradient is large, it is interesting to know how this ”static” decoupling would contribute to the MFMS dynamics. Here, we run simulations of a monochromatic Alfvén wave, using the cHE and pHE stratifications introduced in the previous section. Alfvén waves existing in the low solar atmosphere have been suggested (De Pontieu et al., 2007; McIntosh et al., 2011; Chae et al., 2021), and they may play a vital role in coronal heating (Van Doorsselaere et al., 2020), the FIP effect (Laming, 2004, 2015), etc., which are beyond the scope of this work. We are interested in the fact that the ponderomotive force induced by Alfvén waves drives plasma motions along the magnetic field (Lundin & Guglielmi, 2006), and such motions can be simulated and used as an example to compare the effects of using different stratifications.
An initially constant vertical magnetic field $B_{z}=10$ G is imposed on the stratifications discussed, and the Alfvén wave is driven by periodically changing $B_{x}$ at $z=2.2$ Mm, similar to what was done by Martínez-Sykora et al. (2023). Note that such a driver may not be suitable in the weakly ionized photosphere (Cally, 2023), but works under the present atmospheric condition. The frequency of the driver is 1 Hz, and the amplitude is 1 G. The resulting amplitude of the velocity along the $x$ direction is $\sim 40$ km s-1. Based on this basic setting, five simulations are performed, as listed in Table 1. We mainly focus on comparing the differences between using cHE and pHE stratifications, with or without the NEQ ionization (simulations (c)-(e)). Because of the numerical instability mentioned in the previous section, the pHE stratification is only discussed without including ionization and recombination in the simulation (simulation (d)). We also include two simulations that use non-physical collisional frequencies to examine the weakly and strongly collisional scenarios (simulations (a) & (b)). These two non-physical frequencies are, respectively, significantly lower and higher than the imposed wave frequency. The former should evolve as separated fluids, and the latter should recover the single fluid scenario.
@@
Figure 8: Longitudinal velocities at $t=1$ s. All settings are listed in Table 1. Except simulation (d), which uses the pHE stratification, all the others use the cHE stratification as the initial condition. Simulations (a) and (b) use assigned constant collisional frequencies. More specifically, (a) is weakly collisional and (b) is strongly collisional, respectively having collisional frequencies $5\times 10^{-2}$ Hz and $10^{2}$ Hz. All the other simulations use physical collisional frequencies calculated by Ebysus. Only simulation (e) includes the NEQ ionization/recombination.
@@
The effects of the numerical disturbances appear in the $z$ direction (i.e., the direction of the magnetic field). The Alfvén wave-induced ponderomotive force relates to ${\partial E^{2}_{\perp}}/{\partial z}$, in which $E_{\perp}$ is the electric field perpendicular to the magnetic field (Lundin & Guglielmi, 2006; Martínez-Sykora et al., 2023). Therefore, upward motions (positive longitudinal velocity) are driven by the ponderomotive force, as shown in Figure 8. When it is weakly collisional (panel a), the neutral fluids are decoupled from the ionized fluids, whose motions are driven by the Alfvén wave. Each neutral fluid also moves independently because of the imbalance between gravity and the respective plasma gas pressure, while neither collisions with ionized fluids nor magnetic pressure affect neutral fluids. In contrast, all fluids in the strongly collisional simulation are coupled, behaving as a single-fluid as expected (panel b). More importantly, we can observe that in the strongly collisional simulation, the total velocity is zero before the wave arrives, showing that the cHE stratification is indeed in hydrostatic equilibrium, providing another way to validate the MFMS code – once it is collisionally coupled, the plasma behaves as a single fluid.
@@
Simulations using physical collisional frequencies (Wargnier et al., 2022) show significant decoupling of neutral helium (panels c-e). A similar result was also obtained by Martínez-Gómez et al. (2018), although only initially homogeneous media were considered. Wargnier et al. (2023) also found that neutral helium is more significantly decoupled from other fluids during upper-chromospheric magnetic reconnection. Thus, including helium may be important for modeling the dynamics across the TR. Moreover, when using cHE stratification (panels c and e), neutral helium has a higher longitudinal velocity than pHE (panel d), due to the contribution of the pressure gradient, which is more obvious in Figure 6. The ionization process strengthens the interactions between ionized and neutral fluids, thereby increasing the coupling in the simulation (e). Nevertheless, in general, neutral helium has a smaller longitudinal velocity than the other fluids.
@@
The pHE stratification is not further discussed with ionization and recombination due to the numerical instability. We should, however, note that the ionization and recombination timescales are long in the chromosphere (Carlsson & Stein, 2002), and thus the changes in density and temperature would be less drastic when the pHE stratification is used under chromospheric conditions. However, this also means that the ionization fractions (and thus the collisional interactions) in the corresponding simulations would be largely dependent on the initial stratification. We have shown in Section 2, that heavier elements, whose scale heights are different from that of (ionized or neutral) hydrogen, cannot be properly recovered by the pHE stratification, even if the collisional interactions (or other mixing processes) are sufficient to couple the heavy fluids.